# thanks: Contributed equallythanks: Contributed equally

### v3

@@ III.6 Order-disorder phase transition
We now present how polar order parameter $P=\langle|\mathbf{P}|\rangle_{ss}$ changes with variation of a control parameter $l_{r}/b$ and $l_{t}/b$, shown in Fig. 5. First, for CLF ($l_{t}=0$), phase transitions are shown for different densities of active particles, with area fraction $\phi=\rho\pi b^{2}$ (Fig. 5(a)). With zero $l_{r}$, $P\approx 0$ and there is no flocking (lower bound of Fig. 1(a)). As we increase $l_{r}$, particles flock into a polar ordered state. Whereas with further increase $l_{r}$, disordered phase are observed (see Fig. 1(a)). Transition points shift towards lower $l_{r}/b$ values with increase of density. Then, for the CCF (with fixed $l_{r}=0.55$), polar order parameter $P$ versus $l_{t}/b$ is plotted for different area fractions (for reference, see Fig. 4(b), where the area fraction is fixed at $\phi=0.36$). In our model, the phase transition are highly dependent on the system size and many other parameters such as $\rho$, $v_{0}$, $\chi_{t}$ and $\chi_{r}$. Though of an first-order nature, an extensive study of the exact nature of the transition is left for future work. It needs to be noted that our model includes long-range interactions, and it is computationally expensive to work with such a system.
IV Continuum description
@@
In this section, we write the coarse-grained hydrodynamic equations for the density $\rho(\mathbf{r},t)$ and polarization $\mathbf{p}(\mathbf{r},t)$ fields by appealing to the symmetries and conservation laws of the system [39]. It is also possible to obtain these equations from well-established coarse-graining procedures [17, 38]. The continuum equations for our system can be written as:
$\partial_{t}\rho$ $=-\bm{\nabla}\cdot\left(v_{\text{tr}}(\rho)\,\mathbf{p}-D_{\text{tr}}(\rho)\,\bm{\nabla}\rho\right)$ (14)
$\partial_{t}\mathbf{p}$ $=\Lambda(\rho)\frac{\nabla\rho}{2}+\frac{D_{\text{tr}}[\rho]}{\rho}\Big{[}\nabla\cdot(\mathbf{p}\nabla\rho)\Big{]}-D_{r}\mathbf{p}+D_{t}\nabla^{2}\mathbf{p}$
@@
$v_{\text{tr}}$ $=v_{0}-\chi_{t}\xi_{0}\rho,\quad D_{\text{tr}}=\xi_{t}\rho\chi_{t},\quad\Lambda=\xi_{r}\chi_{r}\rho-v_{\text{tr}}\,\rho$
where the constants $\xi_{0}$, $\xi_{t}$, and $\xi_{r}$ are obtained by fitting the continuum model to the transition line predicted by the particle-based simulation in Fig.1(a) and 3(a). From here on we set $D_{r}$, $D_{t}$ terms to zero in (14) as they are negligibly small as explained in the previous section. In effect, the presence of these terms merely necessitate the formal coarse-graining procedure, but do not play any further role in explaining the instability of the flock.
We now proceed to study the stability of the flocking state. To this end, we linearize (14) about the steady-state density and polarizations ($\rho^{*},\mathbf{p}^{*}$), defining $(\delta\rho,\delta\mathbf{p})=(\rho-\rho^{*},\mathbf{p}-\mathbf{p}^{*})$. We use the following choice of Fourier basis expansion
@@
At the flocking state $(\rho^{*},\mathbf{p}^{*})$ the following conditions hold: $\nabla\rho^{*}\rightarrow 0$ and $\nabla\cdot\mathbf{p}^{*}\rightarrow 0$. We now consider cases of $\chi_{t}=0$ and $\chi_{t}>0$ separately.
IV.1 Case of $\chi_{t}=0$
Here, we can solve equations (14) in real space; we obtain an enslaved dynamics of $\mathbf{\dot{p}}$ with respect to the density gradients:
We see that for the temporal evolution of $\mathbf{p}$ to be stable with respect to the density gradients (feedback response), we require the following.
$\chi_{r}<\frac{v_{0}}{\xi_{r}\rho^{*}}.$ (17)
We see that the fluctuations in both $\rho$ and $\mathbf{p}$ are strongly coupled and that any fluctuations are stabilized in the long-time limit (see Appendix Fig. (7)).
IV.2 General case with $\chi_{t}>0$
Firstly, we may note that solving the $\dot{\mathbf{p}}=0$ in (14) gives $\rho^{*}=\frac{v_{0}}{\xi_{r}\chi_{r}+2\xi_{t}\chi_{t}}$, a generalization of the above inequality. Further, we need to now linearize (14) in the bases of (15) to obtain the stability condition for the flocking state. The full linearization in Fourier space yields the following equation
$\frac{d}{dt}\begin{pmatrix}\delta\tilde{\rho}\\ \delta\tilde{\theta}\end{pmatrix}$ $=\begin{pmatrix}-q^{2}\xi_{t}\chi_{t}\rho^{*}&-v_{\text{tr}}[\rho^{*}]\\ -\frac{q^{2}}{2}B[\rho^{*}]&0\end{pmatrix}\begin{pmatrix}\delta\tilde{\rho}\\ \delta\tilde{\theta}\end{pmatrix}$ (18)
Here $B[\rho^{*}]=(2\xi_{0}\chi_{t}+\xi_{r}\chi_{r})\rho^{*}-v_{0}$. The dispersion relation for $\omega(q)$ then reads
$\omega=\frac{iq^{2}\xi_{r}\chi_{t}\rho^{*}}{2}\Big{[}1\pm\sqrt{1+\frac{2v_{\text{tr}}q^{2}B}{q^{4}\xi_{t}^{2}\chi_{t}^{2}{\rho^{*}}^{2}}}\Big{]}$ (19)
with the condition for long-wavelength instability given by:
$2\xi_{0}^{2}{\rho^{*}}^{2}\chi_{t}^{2}-\xi_{0}\rho^{*}(2v_{0}-\xi_{r}\chi_{r}{\rho^{*}})\chi_{t}-v_{0}(\xi_{r}\chi_{r}\rho^{*}-v_{0})>0$
Using the above, we can solve for $\chi_{t}$. The solution is:
$\chi_{t}^{*}=\frac{v_{0}-\frac{\xi_{r}\chi_{r}\rho^{*}}{2}}{2\xi_{0}{\rho^{*}}}\Big{[}1\pm\sqrt{1-\frac{2(\frac{\xi_{r}\chi_{r}\rho^{*}}{v_{0}}-1)}{(1-\frac{\xi_{r}\chi_{r}\rho^{*}}{2v_{0}})^{2}}}\Big{]}$ (20)
Let us consider the negative solution $\chi_{t}<\chi_{t}^{*}$. Let us further study the limit $\frac{\chi_{r}}{v_{0}}<<1$. Using the definitions for $l_{r}$ and $l_{t}$, the desired inequality is thus
$l_{r}^{*}\gtrsim\frac{2b}{\xi_{r}}\frac{1}{\rho^{*}}+\frac{2.73\xi_{0}b^{3}}{\xi_{r}}l_{t}$ (21)
@@ IV.2 General case with $\chi_{t}>0$
$c_{1}$ $=\frac{2}{\xi_{r}},\quad\quad\beta_{1}=\frac{2.73\xi_{0}b^{3}}{\xi_{r}}$ (22)
We can thus infer the values of $\xi_{r}\approx 2c_{1}\approx 0.026$ and $\xi_{0}\approx 1.6\times 10^{-3}$. We thus conclude that the noiseless limit of a minimal hydrodynamic model captures the destabilization transition of the flock. In addition, such transitions within between order/disorder phases (e.g. Fig. 3(b)) are also captured, although there is no explicit spatial-structure instability encoded in the hydrodynamic model (unlike for instance in [37]).
V Conclusions
@@
To conclude, we present a minimal mechanism for formation of fluid flocks that only requires long-ranged (net) repulsive torques and short-ranged repulsive forces from steric interactions. Although the roles of excluded volume interactions [19, 21, 15, 16], repulsive torques - either short [15] or long [26] ranged - and long-ranged repulsive forces [21, 26] have been studied separately, we have analyzed here the individual contributions of each of those components, and argued that the former two are the minimal requisite components. Our mechanism can summarized as follows: the pair colloids slide together for at least a unit length before deterministically rotating away from each other. This is required to break the forward-backward symmetry at the pair-collision level. Destabilizing of the flock arises due to symmetric collisions at the pair colloidal level, wherein they slide exceedingly short before rotating away upon collision. These results we show to be consistent with a minimal hydrodynamic theory. One may ask whether such a sliding mechanism alone added to cognitive based (agent based) models [10, 11] could reproduce the flocking transition; indeed we will leave this to future work.
@@
The CLF (or the polar liquids) are both generated by, and further destabilized by, the deterministic torques themselves. Notably, the apparent ubiquitous role of noise in destabilizing flocks [4, 10, 21, 26] is also absent in our mechanism (these are negligibly small in our simulation). We note that in the absence of short-ranged excluded volume repulsions, our results would converge to that recently studied in [26] where the need for (long-ranged) repulsive forces is more imminent. Translational repulsion merely renders the flock to acquire a crystalline structure. We also have not studied here the role of particle shape anisotropy [17, 14], which would add additional competing length scales to our analysis. We expect our results here to be directly relevant to a variety of experimental systems. For example, migrating cell collections have been widely reported to exhibit spatial organization resembling a polar fluid [28, 22], while their interaction mechanism has been known to include (among others) long-ranged chemotaxis and avoidance torques [27]. The presented mechanism being very generic, we expect it to also be relevant for interacting colloids of non-chemical origin [25, 38, 37], where the underlying mechanism should also hold.
@@ Acknowledgments
We thank Professor Mike Cates for discussions. AGS acknowledges funding from the DIA Fellowship from the Government of India. SA acknowledges support from the National Postdoctoral Fellowship (SERB File number: PDF/2023/002096) provided by ANRF, Government of India. RS acknowledges support from seed and initiation grants from IIT Madras as well as a Start-up Research Grant from SERB (SERB File Number: SRG/2022/000682), India.
Appendix
Appendix A Simulation details
FIG $L$ dtdt $b$ $v_{0}$ $\kappa$ NN $\chi_{t}$ $\chi_{r}$ $D_{r}$ $D_{t}$
1 $100$ $0.01$ 11 $50$ $175$ $(319,2578)$ 0 $(0,7)$ $10^{-6}$ $10^{-8}$
3.(a) $100$ $0.01$ 11 $50$ $175$ $1024$ $(0,5)$ $(0,10)$ $10^{-6}$ $10^{-8}$
3.(b) $100$ $0.01$ 11 $50$ $175$ $1024$ (1,0)(-1,0) $(0,10)$ $10^{-6}$ $10^{-8}$
3.(d) $186$ $0.01$ 11 $50$ $175$ $4000$ 0 $0.75$ $10^{-6}$ $10^{-8}$
3.(e) $186$ $0.01$ 11 $50$ $175$ $4000$ 33 66 $10^{-6}$ $10^{-8}$
4.(a) $(77,354)$ $0.01$ 11 $50$ $175$ $1600$ 0 $(0,9)$ $10^{-6}$ $10^{-8}$
4.(b) $118$ $0.01$ 11 $50$ $175$ $1600$ $(0,12)$ $(0,32)$ $10^{-6}$ $10^{-8}$
5.(a) $100$ $0.01$ 11 $50$ $175$ $(637,1592)$ 0 $(0,8)$ $10^{-6}$ $10^{-8}$
5.(b) $100$ $0.01$ 11 $50$ $175$ $(637,1592)$ $(0,10)$ $27$ $10^{-6}$ $10^{-8}$
@@ Appendix A Simulation details
Table 1: Parameter values used for respective figures of the paper. Here $L$ is system size, $b$ is the radius of the particle, NN is the number of particles, dtdt is the time-stepping of the Euler-Maruyama integrator. The remaining parameters are defined after Eq.(1). Note that we consider noise to be less dominant to deterministic effects. Indeed, for the results in this paper, the Péclet number $\mathrm{Pe}=v_{0}/(bD_{r})$ is around $10^{9}$.
The update equations is given in Eq. 1 of the main text. The position and orientations of the particles are updated using a forward Euler-Maruyama method. Initially particles are distributed randomly over the two-dimensional space. The initial orientations are also randomly distributed over the range of angles $\pi,+\pi$] (angles are computed with respect to the positive $x$-axis). Periodic boundary conditions are applied on both the $x$-axis and $y$-axis. To compute chemical interactions using the expression of $\mathbf{J}_{i}$, we need to sum over periodic images. We have used a regular summation convention (using minimum image convention) for all results reported here, and checked these with the full Ewald summation convention for selected parameter values. A summary of all of the values of parameters used for the simulation are given in Table (1).
Figure 6: Pair correlation function $g(r,\phi,\theta)$ for the isotropic phase for four different $\theta$ ranges in panel (a). Pair correlation function $g(r,\phi)$ in the band phase for $\chi_{t}=0$ and $\chi_{t}>0$ in (b).
Appendix B Correlation functions and fluctuations

### v4

@@ V.3 Phase transition and finite size effects
We now present how polar order parameter $|\mathbf{P}|$ changes with variation of the control parameter $\Lambda_{r}$ and $\Lambda_{t}$, shown in Fig. 6. First, for CLF ($\Lambda_{t}=0$), phase transitions are shown for different area fraction $\phi=[0.2,0.5]$ with system of size $L=100$ (Fig. 6(a)). With zero $\Lambda_{r}$, $|\mathbf{P}|\approx 0$ and there is no flocking (lower bound of Fig. 2(a)). As we increase $\Lambda_{r}$, particles flock into a polar ordered state. Whereas with further increase $\Lambda_{r}$, disordered phase are observed (see Fig. 2(a)). Transition points shift towards lower $\Lambda_{r}$ values with increase of density. Then we present the finite size effects with different system of sizes $L=[50,400]$ ($N=[239,15279]$) with a constant area fraction $\phi=0.3$ (Fig. 6(b)). The effects of a finite size on the flocking transition can be clearly seen for smaller systems of sizes, whereas the effect is minor for larger system size. Then, for the CCF (with fixed $\Lambda_{r}=0.55$), the polar order parameter $|\mathbf{P}|$ versus $\Lambda_{t}$ is plotted for different area fractions (for reference, see Fig. 5(b), where the area fraction is fixed at $\phi=0.36$). In Fig. 6(d), we study finite-size effects for this case and found the effect on the phase transition is negligible for larger system sizes. It needs to be noted that our model includes long-range interactions, and it is computationally expensive to work with such a system.
Figure 7: (a) Phase diagram in the $(\Lambda_{r},\Lambda_{t})$ plane for the case of $\chi_{t}<0$. The transition line is extrapolated from (3)(a). (b) liquid flock, (c) density band representing the phases of the phase diagram (same color scheme as in Fig (2)). Total number of particles $N=1024$.
VI Effect of long-ranged translational chemo-attraction
To further support the robustness of the mechanism, we report flocks arising from repulsive torques with the addition of long-ranged attractive inter-particle forces, thus $\chi_{t}<0$. We find that collective motion as flocks is maintained for $\Lambda_{t}\in[0,-0.01]$, before transitioning to a disordered state for larger attractive translational interactions; this is shown in Figure 7(a). In this regime, the clumping tendency is balanced by the excluded volume repulsions. This sets the y-intercept of the theoretical prediction in Fig. 7(a) to be of the same order of magnitude to the lower horizontal line in Fig. 2. These are hence liquid flocks (CLF, see Fig. 7(b)). The band formation is also similar to that of the CLF. During density band formation, often less denser waves form in the system, and they move perpendicular to the direction of the density bands. Similar cross-sea phases are also observed in Vicsek-like models [48].
VII Continuum description
@@
To explain the above mechanism, we write top-down hydrodynamic equations for the density $\rho(\mathbf{r},t)$ and polarization $\mathbf{p}(\mathbf{r},t)$ fields by appealing to the symmetries and conservation laws of the system [47]. The continuum equations for our system can be written as:
$\partial_{t}\rho=$ $-\bm{\nabla}\cdot\left(v_{\text{tr}}[\rho]\,\mathbf{p}-D_{\text{tr}}[\rho]\,\bm{\nabla}\rho\right)$ (11a)
$\partial_{t}\mathbf{p}=\,$ $A[\rho]\frac{\nabla\rho}{2}+\frac{D_{\text{tr}}[\rho]}{\rho}\Big[\nabla\cdot(\mathbf{p}\nabla\rho)\Big].$ (11b)
@@
$v_{\text{tr}}[\rho]$ $=v_{0}-\chi_{t}\xi_{0}\rho,\quad D_{\text{tr}}[\rho]=\xi_{t}\rho\chi_{t},$
$A[\rho]$ $=(\xi_{r}\chi_{r}+\xi_{0}\chi_{t})\rho-v_{\text{tr}}[\rho].$
The term $v_{\mathrm{tr}}$ (a coarse-grained speed field) appears in both the cross-coupling terms between density and polarization fields, stabilizing the density field at larger $\chi_{t}$. This effect is in turn countered by $D_{tr}$ (coarse-grained diffusivity), which randomizes the fields at large $\chi_{t}$. The $\chi_{r}$ and $\chi_{t}$ contributions via the $A$ term acts to destabilize the polarization field with respect to density gradients. Given that the transition lines observed are dependent on the area fraction simulated, each of these terms have an explicit density dependence via the constants $\xi_{0}$, $\xi_{t}$, and $\xi_{r}$, which are obtained by fitting the transition line predicted by the particle-based simulation in 3(a). A route to these equations via a systematic coarse-grained procedure is well established [17, 46], though it is not pursued in this work. We note that there are no stochastic contributions appearing in the dynamics of the fields $\rho$ and $\mathbf{p}$ in Eq.(11). We have assumed noise to be negligibly small as explained in the previous section. In effect, the presence of these stochastic terms in the particle-based model is merely necessitated for the formal coarse-graining procedure, but do not play any further role in explaining the instability of the flocking state.
VII.1 Case of $\chi_{t}=0$
First, we may consider the case of only rotational torques. See Fig. 2(a) for phenomenology of this case with $\chi_{t}=0$ from the particle-based model. From (11) - on setting $\chi_{t}=0$ - we obtain an enslaved dynamics of $\mathbf{\dot{p}}$ with respect to the density gradients:
From the above, we see that for the temporal evolution of $\mathbf{p}$ to change the nature of stability with respect to the density gradients (feedback response), we require the following:
$\rho^{*}=\frac{v_{0}}{\chi_{r}\xi_{r}}.$ (13)
We see that the fluctuations in both $\rho$ and $\mathbf{p}$ are strongly coupled and that any fluctuations are stabilized in the long-time limit.
VII.2 General case with $\chi_{t}\neq 0$
For the generic case of both $\chi_{r}\neq 0$ and $\chi_{t}\neq 0$, we may note that solving for $\dot{\mathbf{p}}=0$ in (11), whilst ignoring fluctuations, we obtain:
$\rho^{*}=\frac{v_{0}}{\xi_{r}\chi_{r}+2\xi_{0}\chi_{t}}.$
thus generalizing (13) for $\chi_{t}\neq 0$. In the following, we study the stability of the system about these steady-state values for the density field $\rho^{*}$ and a steady-state value of polar order $\mathbf{p}^{*}$.
VII.3 Stability of the flocking state
To study the stability of the flocking state, we linearize (11) about the steady-state density and polarizations ($\rho^{*},\mathbf{p}^{*}$), defining $(\delta\rho,\delta\mathbf{p})=(\rho-\rho^{*},\mathbf{p}-\mathbf{p}^{*})$. We use the following choice of Fourier basis expansion
@@ VII.3 Stability of the flocking state
For a phase with uniform global polarization, the following conditions will hold: $\nabla\rho^{*}\rightarrow 0$ and $\nabla\cdot\mathbf{p}^{*}\rightarrow 0$.
Eq. (11) can then be linearized in the bases of (14) to obtain the stability condition for the flocking state and subsequently derive the boundary line of Eq. (8), for a generic $\chi_{t}$. The full linearization in Fourier space yields the following:
$\frac{d}{dt}\begin{pmatrix}\delta\tilde{\rho}\\ \\ \delta\tilde{\theta}\end{pmatrix}$ $=\begin{pmatrix}-q^{2}\xi_{t}\chi_{t}\rho^{*}&-v_{\text{tr}}[\rho^{*}]\,\\ \\ -{q^{2}}A[\rho^{*}]/2&0\end{pmatrix}\begin{pmatrix}\delta\tilde{\rho}\\ \\ \delta\tilde{\theta}\end{pmatrix}$ (15)
Here $A[\rho^{*}]=(2\xi_{0}\chi_{t}+\xi_{r}\chi_{r})\rho^{*}-v_{0}$. The dispersion relation for $\omega(q)$ then reads
$\omega=\frac{iq^{2}\xi_{r}\chi_{t}\rho^{*}}{2}\Big[1\pm\sqrt{1+\frac{2v_{\text{tr}}q^{2}A}{q^{4}\xi_{t}^{2}\chi_{t}^{2}{\rho^{*}}^{2}}}\Big]$ (16)
with the condition for linear instability given by:
$2\xi_{0}^{2}{\rho^{*}}^{2}\chi_{t}^{2}-\xi_{0}\rho^{*}(2v_{0}-\xi_{r}\chi_{r}{\rho^{*}})\chi_{t}-v_{0}(\xi_{r}\chi_{r}\rho^{*}-v_{0})<0$
Using the above, we can solve for $\chi_{t}$. For the value:
$\chi_{t}^{\pm}=\frac{v_{0}-\frac{\xi_{r}\chi_{r}\rho^{*}}{2}}{2\xi_{0}{\rho^{*}}}\Big[1\pm\sqrt{1+\frac{2(\frac{\xi_{r}\chi_{r}\rho^{*}}{v_{0}}-1)}{(1-\frac{\xi_{r}\chi_{r}\rho^{*}}{2v_{0}})^{2}}}\Big]$ (17)
we can consider the solution $\chi_{t}>\chi_{t}^{-}$. Let us further study the limit $\frac{\chi_{r}}{v_{0}}<<1$, which is corresponds to parameter ranges used in our simulations (see Table 1).
Using the definitions for $\Lambda_{r}$ and $\Lambda_{t}$ - see Eq.(5) - the condition for instability of the modes $\omega(q)$ is given by
$\Lambda_{r}^{*}\gtrsim\frac{2}{3b^{3}\xi_{r}\rho^{*}}-\frac{1.67b\xi_{0}}{\xi_{r}}\Lambda_{t}$ (18)
@@
$\alpha$ $=\frac{2}{3b^{3}\xi_{r}},\quad\quad\beta=-\frac{1.67b\xi_{0}}{\xi_{r}}$ (19)
We can thus infer the values of $\xi_{r}\approx 2/(3b^{3}\alpha)\approx 51.3$ and $\xi_{0}\approx-5.13/b$. The hydrodynamic instability of the flocking phase thus captures the transition lines between ordered and disordered phase in Figures 2(a) and 3(a). Though Eq. (18) captures the transition for a generic crystalline flock (generic $\chi_{t}$), we note that there is no explicit spatial-structure instability encoded in the hydrodynamic model [44], thus if one simulates Eq. (11) one obtains effectively fluidic flocks at a continuum level. Nevertheless, we conclude that the noiseless limit of a minimal hydrodynamic model captures the destabilization transition of the flock, with exact details of the spatial structure of the system likely requiring to go beyond the conventional molecular chaos hypothesis [16, 17].
VIII Summary and discussion
@@
To conclude, we present a minimal mechanism for formation of fluid flocks that only requires long-ranged (net) repulsive torques and short-ranged repulsive forces from steric interactions. Although the roles of excluded volume interactions [20, 31, 22, 16], repulsive torques - either short [22] or long [26] ranged - and long-ranged repulsive forces [31, 26] have been studied separately, we have specified here the individual contributions of each of those components, and argued that the former two are the minimal requisite components. The mechanism can summarized as follows: the pair colloids slide together for at least a unit length before deterministically rotating away from each other. This is required to break the forward-backward symmetry at the pair-collision level. Destabilizing of the flock arises due to symmetric collisions at the pair colloidal level, wherein they slide exceedingly short before rotating away upon collision. These results we show to be consistent with the noiseless limit of a minimal hydrodynamic theory. One may ask whether such a sliding mechanism alone added to cognitive based (agent based) models [13, 14] could reproduce the flocking transition; indeed we will leave this to future work.
@@
It is important to note that the CLF and CCF phases are both generated by, and further destabilized by, the deterministic torques themselves, independent of the contribution of noise. Though we have explicit noise terms in the updates in Eq.(1) for the sake of completion, they are set to be negligibly small (see Table 1 in Appendix). We note that we have repeated all the simulations presented in Fig. (2) with $D_{r}=D_{t}=0$, and have found the results to be identical. Thus, our results on flocking of phoretically interacting active particles and their destabilization are distinct from conventional picture of noise-induced inhibition of flock formation [6, 13, 31, 26]. We note that in the absence of short-ranged excluded volume repulsions, our results on the formation of CLF would converge to that recently studied in [26] where the need for (long-ranged) repulsive forces is more imminent. Translational repulsion merely renders the flock to acquire a crystalline structure. We also find flocks for (a narrow range of) long-ranged attractive forces, if of the order of less than the short-ranged repulsion. We have not studied here the role of particle shape anisotropy [17, 49], which would add additional competing length scales to our analysis. We have also not studied the role of hydrodynamic interactions between the particles [3, 50, 51]. We expect our results here to be directly relevant to a variety of experimental systems. For example, migrating cell collections have been widely reported to exhibit spatial organization resembling a polar fluid [33, 18], while their interaction mechanism has been known to include (among others) long-ranged chemotaxis and avoidance torques [32]. The presented mechanism being very generic, we expect it to also be relevant for interacting colloids with non-diffusiophoretic fields [25, 46, 44], where the underlying mechanism should also hold.
@@ Acknowledgments
We thank Professor Mike Cates for discussions. We also extend our gratitude to two anonymous reviewers of the manuscript for their feedback and constructive criticism, which led to an improvement in the presentation of our results. AGS acknowledges funding from the DIA Fellowship from the Government of India. SA acknowledges support from the National Postdoctoral Fellowship (SERB File number: PDF/2023/002096) provided by ANRF, Government of India. RS acknowledges support from seed and initiation grants from IIT Madras as well as a Start-up Research Grant from SERB (SERB File Number: SRG/2022/000682), India.
Appendices
.1 Table of parameters
A table of all the parameters used for the generated figures is presented in Table 1.
Figure no. $L$ dtdt $b$ $v_{0}$ $\kappa$ NN $\phi$ $\chi_{t}$ $\Lambda_{t}$ $\chi_{r}$ $\Lambda_{r}$ $D_{r}$ $D_{t}$
2 $100$ $0.01$ 11 $50$ $175$ $(319,2578)$ $(0.1,0.8)$ 0 0 $(0,7)$ $(0.0,0.14)$ $10^{-6}$ $10^{-8}$
3.(a) $100$ $0.01$ 11 $50$ $175$ $1024$ $0.32$ $(0,5)$ $(0.0,0.1)$ $(0,10)$ $(0.0,0.2)$ $10^{-6}$ $10^{-8}$
4.(a),(c) $186$ $0.01$ 11 $50$ $175$ $4000$ $0.36$ 0 0 $0.75$ $0.015$ $10^{-6}$ $10^{-8}$
4.(b),(c) $186$ $0.01$ 11 $50$ $175$ $4000$ $0.36$ 33 $0.06$ 66 $0.12$ $10^{-6}$ $10^{-8}$
5.(a) $(77,354)$ $0.01$ 11 $50$ $175$ $1600$ $(0.04,0.84)$ 0 0 $(0,9)$ $(0,0.18)$ $10^{-6}$ $10^{-8}$
5.(b) $118$ $0.01$ 11 $50$ $175$ $1600$ $0.36$ $(0,12)$ $(0,0.6)$ $(0,32)$ $(0,0.64)$ $10^{-6}$ $10^{-8}$
6.(a) $100$ $0.01$ 11 $50$ $175$ $(637,1592)$ $(0.2,0.5)$ 0 0 $(0,8)$ $(0.0,0.16)$ $10^{-6}$ $10^{-8}$
6.(c) $100$ $0.01$ 11 $50$ $175$ $(637,1592)$ $(0.2,0.5)$ $(0,10)$ $(0.0,0.2)$ $27$ $0.54$ $10^{-6}$ $10^{-8}$
7.(a) $100$ $0.01$ 11 $50$ $175$ $1024$ $0.32$ (1,0)(-1,0) (0.02,0.0)(-0.02,0.0) $(0,10)$ $(0.0,0.2)$ $10^{-6}$ $10^{-8}$
@@
Table 1: Parameter values used for respective figures of the paper. Here $L$ is system size, $b$ is the radius of the particle, NN is the number of particles, dtdt is the time-stepping of the Euler-Maruyama integrator. The remaining parameters are defined after Eq.(1). We have kept $\frac{c_{0}}{4\pi D_{c}}=1$ in all the simulations. Note that we consider noise to be less dominant to deterministic effects. Indeed, for the results in this paper, the Péclet number $\mathrm{Pe}=v_{0}/(bD_{r})$ is around $10^{9}$.
Appendix A Pair correlation functions